Tohban Report 2014-02-12

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 05 Fed 2014warning.pngThe date "05 Fed 2014" was not understood.
End Date: 12 Feb 2014
Tohban: Nicole Duncan
Tohban email:
Next Tohban: Hazel Bain
List all reports


Solar Activity

Solar activity dropped off early/mid way through the week, but picked back up by the end of the week. We are currently under a Major Flare watch. The most active region on disk remains AR 11974 with a beta-gamma-delta configuration. Its in disk center and its expected to maintain activity over the coming week. Stereo B doesn't show any complex ARs to rotate on disk in the next week. Maybe by the end of the week, or early next week we should see another AR rotate on disk.

How many GOES flares occurred?

 Flares above B, C, M, X class were      0    52    9     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      0    40    7     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0    18     4     0

There were RHESSI flares/GOES flares 619 / 61 over the time range 04-Feb-14 11-Feb14

Memory Management

With added passes and a quieting of the sun, the SSR emptied early in this week. It is increasing over the past day. This morning started .

Spacecraft Status

Nominal. The cold plate 1 and 2 are at about 114.9 K and 113.5 K respectively, down from last week. Cryocooler performance normal.

Data Gaps

GAPS IN APP_ID = 1 (VC1-SOH) WITH PACKET RATE LT 30 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2014-02-10T17:55:00.000 -- 2014-02-10T18:05:00.000 600.00000

GAPS IN APP_ID = 154 (VC1-PMTRAS) WITH PACKET RATE LT 4 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2014-02-10T18:00:00.000 -- 2014-02-10T18:10:00.000 600.00000

GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 GAP START TIME GAP END TIME GAP (SEC) 2014-02-11T18:40:00.000 -- 2014-02-12T00:00:00.000 19200.000

Detector issues

On Feb 5th commanding was done to increase D6's front slow threshold at 2014-036-20:27:13 UTC

        was     now

6F: 0x0C 0x1A (to about 6 keV)

Detector 2 rear resets continue to be seen this week.

Other notes

Fronts and rears should have been turned off during eclipse. But it seems like the fronts are on at night (they're not supposed to be), which would increase our SSR. Today this will be checked, and if necessary, commanding will be done.

Spacecraft Management

No operations this week

Decimation Active/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) Enabled 4 minutes before and after eclipse
Night time data (rears) Disabled, during entire eclipse
Require extra passes? No?
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? None
Personal tools